OpenStudy (esshotwired):

First person gets a medal! Three brothers live in a farm. They agreed to buy new seeds: Adam and Ben would go and Charlie stayed to protect fields. Ben bought 75 sacks of wheat in the market whereas Adam bought 45 sacks. At home, they split the sacks equally. Charlie had paid 1400 dollars for the wheat. How much dollars did Ben and Adam get of the sum, considering equal split of the sacks?

OpenStudy (esshotwired):

It is a riddle and you have to solve it correctly

OpenStudy (esshotwired):

if you HAVE to, you can look it up online, but if you know it you will have a better chance at getting the medal.

OpenStudy (esshotwired):

no you actually have to find how much ben gets and adam gets individually

OpenStudy (anonymous):

75+45=120 120/3=40 1400/40=35ea Adam: 35(45-40)=$175 Ben: 35(75-40)=$1225

OpenStudy (esshotwired):

Yes!!